A r = (X r ,Y r ), r=1,2,3………,n is a set of n points. P 1 divides A 1 A 2 in the ratio 1:1, P 2 divides P 1 A 3 in th ratio 1:2, P 3 divides P 2 A 4 in the ratio 1:3 and so on . P n-1 divides P n-2 A n in the ratio 1:n-1, Find The Co-ordinates of P n-1 ?


Ar = (Xr,Yr), r=1,2,3………,n is a set of n points. P1 divides A1A2 in the ratio 1:1, P2 divides P1A3 in th ratio 1:2, P3 divides P2A4 in the ratio 1:3 and so on . Pn-1 divides Pn-2An in the ratio 1:n-1, Find The Co-ordinates of Pn-1?


Grade:11

1 Answers

Askiitians_Expert Yagyadutt
13 years ago

Hii vinay !

 

This is very nice question ....Try to understand this -

 

 

A1     P1         A2            P2           A3             P3            A4             P4               A5                  .........Pn-1.......  An

 

(x1,y1)          (x2,y2)                 (x3,y3)                                                                                                      (xn,yn) 

 

P1   divides A1 A2 in 1:1

 

so P1 = [(x1 +x2)/2 ,(y1+y2)/2 ]

 

P2 divides P1A3 is 1 : 2 ...

 

so P2 =  [2P1 + A3 ]/3   =   [ (x1+x2+x3)/3 , (y1+y2+y3)/3 ]

 

Similarly in P3  you will get   [ (x1+x2+x3+x4)/4 , (y1 + y2 + y3 +y4)/4]

 

And thus Pn-1 = [ (x1+x2+x3........+xn)/n , )y1+y2+y3........+yn)/n ]

 

 

Note : simply use the section formula which means if any point (h,k) divide (p,q) and (r,s) in ration a:1

 

(p.q)...............a.........(h,k)...................1............(r,s)

 

Then  h  = (a.r + p )/(a+1)   k = ( as + q)/(a+1) 

 

 

With regards

 

yagya

askiitians_expert